i need to find the value of x in terms of a and b in the simplest form assumeing (3a = -2b)for the follwoign equation
3a(3a-x)=2b(x+2b)
thanks

Answer by stanbon(75887)

find the value of x in terms of a and b in the simplest form assumeing (3a = -2b)for the follwoign equation
3a(3a-x)=2b(x+2b)
9a^2 - 3ax = 2bx + 4b^2
----
2bx+3ax = 9a^2-4b^2
----
(2b+3a)x = 9a^2-4b^2
x = [(3a-2b)(3a+2b)]/(3a+2b)
x = 3a-2b
-----
If 3a = -2b you get:
---
x = -4b
==============
Cheers,
Stan H.